OpenStudy (anonymous):

Evaluate the integral using integration by parts where possible x^5 lnx dx

OpenStudy (uri):

u = lnx => du = 1/x dv = x^5 => v = x^6/(6)

OpenStudy (uri):

then use IBP formula. Sub it in.

OpenStudy (anonymous):

I got \[1/6 x^6lnx-1/4x^2+C\] can you tell me what I did wrong?

OpenStudy (uri):

Show me your working out.